Hubble Space Telescope Cycle 5 General Observer Proposal

Imaging of Two Dwarf Galaxies in Tidal Tails

Principal Investigator: John Hibbard
Institution: University of Hawaii

Institute for Astronomy 2680 Woodlawn Drive Honolulu, HI 96822

Telephone: 808-956-8312
Electronic mail: hibbard@galileo.ifa.hawaii.edu
ESA member: No
Scientific category: Galaxies & Clusters
Instruments: WFPC2
Cycle 5 primary orbits: 11
Cycle 5 parallel orbits: 0

Abstract

We propose to image the two nearest dwarf galaxy candidates that are entrained in tidal tails. NGC 3077 is a member of the M81 group, and recent models suggest that it is interacting with, but has not formed from, the HI surrounding M81. In contrast, a dwarf at the SW end of the NGC4038/9 interaction tail appears to have formed out of gas ejected from that merger. The latter process is proposed to be a significant mechanism for the formation of dwarf spheroidal galaxies (cf. Schweizer, 1978; Barnes & Hernquist, 1992). We will determine the ages, luminosity functions, and spatial distributions of OB associations in these systems and also search for older (> 1 Gyr) red AGB stars. We hope to illuminate the star formation history of these dwarfs and constrain how much of their mass (particularly the NGC4038/9 system) has formed after the merger event. We may also shed light on the long term survivability of the 4038/9 dwarf as a stellar system.

John Hibbard

Imaging of Two Dwarf Galaxies in Tidal Tails

Investigator  Institution  Country
PI:  John Hibbard  University of Hawaii  USA/HI
CoI: R. Michael Rich  Columbia University  USA/NY
Total number of investigators: 2
Number of ESA investigators: 0 (indicated by ^{∗} after name)

Observing Summary:  Total
Target RA DEC V Configuration,mode,spectral elements orbits Flags

NGC 3077 10 03 68 44 10.7 WFPC2 IMAGE

F336W,F439W,F555W,F814W 1

CVZ,DUP

NGC 4038/9-DWARF 12 01 −19 01 17.3 WFPC2 IMAGE

F336W,F439W,F555W,F814W

10

Grand total orbit request  11

Scientific Justification

1 Introduction

It was almost four decades ago that Zwicky first suggested that dwarf galaxies might condense out of the luminous filements protruding from and between galaxies (Zwicky 1956). Schweizer followed up this suggestion in 1978 with a detailed spectroscopic study of several clumps imbedded within these tidal tails (Schweizer 1978). Central to this work were detailed observations of the classic merger NGC 4038/9. Using the CTIO 4m reflector, Schweizer clearly discovered a condensation at the Southwest end of one of the tails, referring to it as a “dwarf stellar system” (Fig. 1). Condensations in the HI tails of NGC 7252 (Hibbard et al. 1993) and theoretical work by Barnes & Hernquist (1992) have further spurred the idea that dwarf galaxies might condense out of the expanding tidal debris associated with galactic mergers. Imaging and spectroscopy of the Southwest knot in the NGC 4038/9 system by Mirabel, Dutton, & Lutz (1992) confirm in detail Schweizer’s claim that numerous HII regions are present, and that consequently massive stars are in the act of forming. However, for a dwarf spheroidal system to stand alone, it must be self-gravitating. The current formation of stars and substantial amounts of HI (Mahoney, van der Hulst, & Burke, 1994) is not sufficient to assure that the condensation will actually survive as an independent galaxy many Gyr after the conclusion of the merger.

We are proposing to explore the two nearest dwarf systems imbedded within neutral hydrogen “tails”. In the case of the NGC 4038/9 knot, there is much evidence (e.g. Schweizer’s photometry) that the colors are extremely blue and that gas is forming into stars for perhaps the first time. The luminosity function and stellar population of this system will be of great interest, as will the presence or absence of older red stars. As a contrast, we propose to image NGC 3077, the dwarf companion of M81. It too is entrained within a gaseous tidal tail (van der Hulst 1979, Yun 1992), but theoretical models of the merger (Yun, 1994) point toward NGC 3077 being an interloper, not having formed from tidal debris. Both galaxies are near enough that deep imaging will resolve stars and could reveal a 1 Gyr old AGB population if it is present. These data can be used to explore star formation in this most unusual context and may potentially reveal the structure of each of the galaxies, as delineated by older stars.

2 Significance to Astronomy

While still a subject of some debate, it is clear that galaxy interactions are an important part of the evolution of galaxies. It is possible that a fair fraction of field ellipticals are merged spirals (Toomre & Toomre 1972, Schweizer & Seitzer 1992). Material from merger tails is potentially responsible for quasar absorption line systems (Morris & van den Bergh 1994), as mergers are easily capable of distributing material over 100 kpc or more (Hibbard 1994). As with any process (and its details) with increasing widespread popularity, there is a temptation to evoke galactic merging to explain a wide variety of phenomena, perhaps even to a fault. The idea that dwarf galaxies may form from merger ejecta is potentially quite Figure 1: Dwarf system in the merging galaxy NGC 4038/9. The small image illustrates the entire merger; the enlargement is a 3X3 arcmin square centered on the knot and dwarf at the lower right (SW) end of the tail. (North at top, East to the left). Smaller square is the size of the PC. A WFPC2 field would approximately cover the area shown, resolving the stellar population. The dense knots are extremely high excitation HII regions studied by Mirabel, Dottori & Lutz, 1992, A&A, 256, L19. The brightest blue stars would be resolved at V = 21.5 − 22, for (m = M) = 31.5. interesting. However, optimistic assertions in the literature (cf. Mirabel et al. 1992) fail to show that the potential dwarfs are self-gravitating and thus would be capable of surviving independently long after the merger signatures are gone. Imaging of the resolved stellar content is likely going to reveal if the starlight is sufficiently concentrated to bind the new system and its gas.

The star formation history, spatial distributions of older stars, OB associations, etc. will constrain the fraction of stars formed out of tail gas. Beyond the issue of the merger itself is the interest in exploring star formation in such an unusual environment. Mirabel et al. (1992) find Hβ line-strengths above that of 30 Doradus and comparable to those of the unusual blue compact galaxy II Zw 40; stars as young as 2 × 10^{6}yr are predicted (at M_{V} = −10, these stars will appear at V ≈ 21.5). For the NGC 4038/9 condensation and NGC 3077, we want to know if the luminosity function follows Freedman’s (1985) slope established for the Local Group. It is interesting that the conditions prevailing in the dwarf may have been similar to those encountered in galaxy formation.

NGC 3077 is a peculiar galaxy in its own right, but should have a substantial older population. It will nonetheless be interesting to explore the spatial distribution of any recent star formation, given this galaxy’s presence in the M81 neutral Hydrogen interaction stream. Our exposures should reach to V=27, or M_{V} = −1, easily deep enough to reveal a red AGB. Should such a population be absent it would be convincing evidence that this system condensed out of the merger gas rather than entering the merger from outside.

These observations will help to answer an interesting observational question. They will also contribute valuable observational material to help constrain current theoretical work. In particular, if such condensations are assembled from the growth of pre-existing irregularities within the expanding outer disk material, as in the models of Barnes & Hernquist (1992), they should show a substantial older population. If, on the other hand, they form wholly from a jeans instability within the expanding gaseous tails, as suggested by Elmegreen et al. (1993), they should have relatively few old stars.

3 Proposed Observations

We will image the region of the 4038/9 tail that has the greatest activity, the region where Mirabel et al. 1992 postulate a dwarf may be forming right now. We propose to reach at least V=28 in the NGC 4038/9 dwarf; a recession speed of 1600 km/sec (≈ 20 Mpc) corresponds to a modulus of 31.5. We can potentially reach to M_{V} ≈ −3.5 and could map association ages up to 28 Myr (turnoff mass of 9M_{⊙}). This would be an effort comparable to Wilson’s (1991) study of OB associations in M33. Note that while the distance to 4038/9 is uncertain by 0.5 mag, it will still be possible to explore ages up to ≈ 25 Myr. We will easily be able to see the extent of current and recent star formation.

It is interesting that these observations might help to constrain the distance of the whole interaction. Given that the Hβ equivalent width is consistent with ionization from stars with < 2 Myr (or M_{V} = −10) it may be possible to constrain the distance to this galaxy using the brightness of the young main sequence stars.

A population old enough to have an asymptotic giant branch should have red stars visible at I ≈ 26.5. These giants should trace the mass as represented by the old stellar population. The occurrence of even a trace of such stars would be strong evidence that the bulk of the “dwarf” consists of tail stars that became self-gravitating, drawing more stars and gas in to form valid dwarf systems.

The older population and its spatial distribution will be easily revealed in NGC 3077. In only 2 orbits, one can reach as deep as is possible for NGC 4038/9. Using a modulus of 27.8 (Freedman et al. 1994) one reaches to M_{V} = −1, easily deep enough to reach both the young population and intermediate-age stars.

One of the goals in constructing the Hubble Space Telescope was the hope that we would be able to conduct detailed studies of objects and problems limited by spatial resololution. Here we have the opportunity to study the detailed content of two galxies, one which may be recently formed and portray a process of galaxy formation. The object in question is unlike anything that has been studied in our vicinity, and would be inaccessible without the spatial resolution offered by HST.

4 REFERENCES

Barnes, J.E., Hernquist, L., 1992. Nature, 360, 715. Elmegreen, B., Kaufmann, M., Thomasson, M., 1993, ApJ, 412, 90. Freedman, W.L. 1985 ApJ 299, 74. Freedman, W.L. et. al. 1994, ???. Hibbard, J.E., 1995, Ph.D. Thesis, Columbia University. Hibbard, J.E., Guhathakurta, P., van Gorkom, J.H., Schweizer, S., 1994, AJ, 107, 67. van

der Hulst, J.M., 1979, A&A, 155, 151. Mahoney, J.H., van der Hulst, J.M., Burke, B.F., 1994, in preparation. Mirabel, I.F., Lutz, D., Maza, J., 1991, A&A, 243, 367. Morris, S.L., van den Bergh, S., 1994, ApJ, 427,696. Schweizer, F., 1978. “The Structure and Properties of Nearby Galaxies”, p. 279. IAU Symp.

No. 77. E.M. Berkhuijsen and R. Wielebinski, eds. Schweizer, F., Seitzer, P., 1992, AJ, 104, 3. Toomre, A., Toomre, J., 1972, ApJ, 178, 623. Wilson, 1991, AJ, 101, 1663. Yun, 1992, thesis, Caltech. Yun, M.S., et al. 1994, Nature, submitted. Zwicky, F., 1956, Ergebnisse der Exakten Naturwissenschaften, 29, 344.

Description of the Observations

Two sets of observations are proposed. (1) “light” U, B exposures; 1/2 orbit each for 3077 and 1 orbit each for 4038/9. These are used to reach the brightest OB stars and measure the reddening. Deep V, I exposures will obtain the luminosity function and spatial distribution of the young stars, and will reach deep enough to measure 1 Gyr old AGB stars, should they be present. A galaxy formed entirely from a recent merger should have no stars on the AGB.

We consider the deep integration of 4 orbits in each of V (F555W) and I (F814W) for the NGC 4038/9 dwarf. If each orbit is to be CR-split, then no integration is longer than about 1200-1500 sec. We modify Eqn. 6.2 to read that (S/N)^{2} = I^{2}/(B_{tot} + I) where B_{tot} is either 9 (WF) or 25 (PC) times the background rate per pixel. We use the appropriate QT integral for the filter and instructions on p. 58 to find that the sky (ecliptic plane) will beat the read noise in 1000 sec (V=22.3 + 5 gives 87 c/pix). Formally, one should reach V ≈ 27 with S/N = 3 in this time. In practice, the cycle 4 data of Rich and Mighell on Leo II (at −88^{∘}C) reach V = 27 with errors of 0.5 mag in 4800sec. We expect that with great care, we can reach V = 28 with 0.3 mag errors in the 10,000 sec available in 4 orbits. This will reach to M_{V} = −3.5, about as deep as Wilson’s (1991, AJ, 101, 1663) study of OB associations in M33. The presence of red stars at M_{I} = −4.5 would prove an old population is present. We must reach I = 26.5 to solidly detect these stars. In a 1000 sec exposure, the background/pixel is 93 cts, and 112 cts is detected from the star. S/N = 3.6 in this time, and one may hope to reach I = 28 in the full 4 orbits. These estimates are more conservative than the WFPC2 handbook, but agree with what we can do with our deep V,I images of the Leo II dwarf and have been confirmed (at the time of writing) by another group with deep data.

For NGC 3077, (m − M) = 27.8. In 2 orbits we get ≈ 5000 sec in each color due to the efficiency of the CVZ. We estimate that we could reach V ≈ 28 and I ≈ 27, which would definitely permit us to detect any older stars present.

Special Requirements

NONE.

Supporting/coordinated Observations

Hibbard plans a deep neutral hydrogen map with the VLA. Deep infrared imaging is also planned for both NGC 3077 and the 4038/9 dwarf, using telescopes on Mauna Kea. The IR maps will be compared with the young star distribution to estimate the ratio of new to old stellar population.

Related HST Programs

J. Hibbard: No HST data.

R. M. Rich (RMR)

RMR GO 2735 (Cycle 1)/5019(Cycle 3) “A Survey of the Giant Branch in the Nuclear Bulge of M31” R. M. Rich, PI All data in this program have been reduced and a 3rd journal paper is in preparation (in addition to those listed below).

“Stellar Photometry in the Inner Bulge of M31 using the Hubble Space Telescope Wide Field

Camera” ApJ in press (20 Jan 1995) “Stellar Photometry in the Inner Bulges of M31 and M33” Mighell, K. & Rich, R.M. 1994

in The Local Group (ESO: Garching) in press. “The Morphology of M31’s Nucleus from 0.5 to 3.5 μm” Pahre, M.A., Mould, J.R., Mathews,

K., Mighell, K., & Rich, R.M. PASP submitted October 1994 4 abstracts at June, 1993 AAS meeting (Berkeley) RMR GO 3804 (Cycle 3) “Imaging of UV-bright stars in Metal Rich Galactic globular

clusters” Rich, R.M. (PI), Liebert, J., King, I. Data reduced. An anomalous hot HB has been discovered in the metal rich cluster NGC6388 and a paper is in preparation. FOC exposures failed to reach sufficiently deep to discover UV-bright stars in the 4 other clusters observed, but those data can be used to study the stellar distribution.

RMR GO 4546 (Cycle 3) Imaging of UV-bright stars in the Galactic Bulge; Rich, R.M. (PI), Liebert, J., King, I., Renzini, A., Greggio, L. Observations seriously compromised by failure of FOC/48. 64 frame scan in F140W with 20 stars detected; no comparable optical data to identify and remove red leak stars; analysis is in progress.

RMR GO 5386 (Cycle 4 complete) “Main Sequence Photometry of the Leo II Dwarf Spheroidal Galaxy”; Mighell, K. (PI), Rich, R.M. Data reduced; main sequence turnoff detected for the first time, along with a curious blue straggler population. We expect to submit a paper by the end of 1994.

RMR GO 5475 Shara, M. (PI), Rich, R.M., Mighell, K., Fall, M. “A Survey for Blue Stragglers in Magellanic Globular Clusters” Cycle 4 snapshot; most data obtained; Preliminary reductions have been done; these data were taken early in Cycle 4 without CR-split and have serious problems with warm pixels; recalibration is necessary and reductions are complicated. Expect first papers by end of 1994.

RMR GO 5464 (Cycle 4; no data yet) “A Survey of the AGB in Local Group Bulge Popu-

lations” Rich, R.M., Mighell, K., Freedman, W. RMR (Based on HST Data) “Population Gradients and Supra-Horizontal Branch Stars in

the Galactic Center Globular Cluster NGC 6522” Shara, M., Drissen, L., Rich, R.M., Paresce, F., King, I., & Meylan, G. Submission to PASP expected in October, 1994.

“Stellar Populations Near the Nucleus of M33” Mighell, K.J. & Rich, R.M. Submission to AJ expected in September, 1994.

Justify Duplications

A Search of the Archive finds one image (Cycle 2) of NGC 3077 in FOC/96 (600 sec) F220W (only color). SNAP 4804/MAOZ. Our new multicolor data will be much deeper and will permit construction of a color-magnitude diagram.